Preferred

stockholder expected

return​)

You own

100

shares of Shapard Resources preferred​ stock, which currently sells for

​$37

per share and pays annual dividends of

​$5.25

per share.

a. What is your expected​ return?

b. If you require a return of

11

​percent, given the current​ price, should you sell or buy more​ stock?

a. Your expected return is

nothing

percent.  ​(Round to two decimal​ places.)b. If you require a return of

11

​percent, the value of the stock for you is

​$ .

​(Round to the nearest​ cent.)Because the expected rate of return is

less than

greater than

your required rate of return or the intrinsic value of because the current market price is

less than

greater than

​$47.73​,

the Shapard Resources preferred stock is

overvalued

undervalued

and you should

sell

buy

the stock.  ​(Select from the​ drop-down menus.)
Answer #1

a. What is your expected​ return?

=5.25/37=14.189%


b. If you require a return of 11 percent, given the current​ price, should you sell or buy more​ stock?

Value of the stock=5.25/11%=47.72727273


Because the expected rate of return is greater than your required rate of return or the intrinsic value of 47.73 because the current market price is less than $47.73​, the Shapard Resources preferred stock is undervalued and you should buy the stock.